OpenStudy (anonymous):

Given that y=1 when x=0, solve the equation dy/dx= y(4-y) Obtaining an expression for y in terms of x

OpenStudy (anonymous):

the first step to solving an initial value differential equation is that you must separate the variable ;)

OpenStudy (shubhamsrg):

re-arrange : dx = dy/(y(4-y)) integrate both sides..a constant will come up .. use y=1 for x= 0 and find constant..thus you get the soln..

OpenStudy (anonymous):

How to integrate both sides?

OpenStudy (anonymous):

∫dx=∫dy/(y(4-y))

OpenStudy (shubhamsrg):

1/(4y - y^2 ) = (4-2y)/(4y - y^2) + (-3 + 2y)/(4y-y^2) the second one can be done with more than 1 methods..use partial differentiation.. in first one..note that d/dy(denominator) = numerator.. hope this helps..

OpenStudy (anonymous):

Partial fractions \[\frac{1}{y(4-y)}=1/4(\frac{1}{y}+\frac{1}{4-y})\]

OpenStudy (anonymous):

x+C=∫dy/(4-(y-2)^2); that is obtained by completing the square

OpenStudy (anonymous):

Integrate \[x=\frac{1}{4}[\ln(y)-\ln(4-y)]+c\] Put x,y=(0,1)to find c

OpenStudy (anonymous):

Is that all?

OpenStudy (anonymous):

Pretty much

OpenStudy (anonymous):

Is this the same? y = 4 /(3 e^−4x + 1)?